OpenStudy (anonymous):

point slope : parallel to 3x-6y=5 , containing (-2,4)

OpenStudy (anonymous):

parallel to 3x-6y=5 means they have the same slope, to find this just divide the number in front of the x by the negative of the number in front of the y, so m=3/-(-6) m=1/2 point slope is of the form y-y1=m(x-x1) so plugging in y1 and x1 from what you have gives you y-4=1/2*(x+2)